40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15445     Entry time: Wed Jul 1 12:50:40 2020
Author: gautam 
Type: Update 
Category: PEM 
Subject: MC1 accelerometers plugged in 

I re-connected the 3 accelerometers located near the MC1/MC3 chamber. It was a bit tedious to get the cabling sorted - I estimate the cable is ~80m long, and the excess length had to be wound around a spool (see Attachment #1), which wasn't really a 1 person job. It's neat-ish for now, but I'm not entirely satisfied. I think we should get shorter cables (~20m), and also mount the pre-amp/power units in a rack instead of leaving it on the floor. The pre-amp settings are x100 for all three channels. The MC2 channels are powered, but are unconnected to the seismometers - it was too tedious to unroll the other spool yesterday. Apart from this, the cable for the "Z" channel had to be re-seated in the strain relief clamp.

I did not enable any of the CDS filters that convert the raw signal into physical units, so for now, these channels are just recording raw counts.

Update 7pm: the spectra in the current config are here - not sure what to make of the MC2_Z channel appearing to show lower noise?

Update July 13 2020 430pm: This afternoon, I hooked up the MC2 accelerometer channels too...

Attachment 1: IMG_8617.JPG  2.718 MB  Uploaded Wed Jul 1 13:54:06 2020  | Hide | Hide all
IMG_8617.JPG
Attachment 2: IMG_8616.JPG  2.987 MB  Uploaded Wed Jul 1 13:55:44 2020  | Hide | Hide all
IMG_8616.JPG
ELOG V3.1.3-